# Local Attention Mechanism: Boosting the Transformer Architecture for Long-Sequence Time Series Forecasting

### v2

@@ 2.4 The Encoder-Decoder architecture
3 Local Attention Mechanism (LAM)
In this section we introduce our proposal, the LAM mechanism, and an efficient algorithm, based on expressing LAM in tensor algebra, that leads to $\Theta(n\log n)$ memory and time complexity. This algorithm has been implemented in stock Pytorch and TensorFlow without the need for any low level CUDA code.
The high-level fundamental behind LAM is exploiting locality in order to re-define the attention mechanism. Previous studies on time series have empirically shown that the rows of the attention scores matrix behave as long tail distributions [28, 22]. This behaviour is present at any layer independently of the depth. For example, the coefficients $a_{n-1,0},a_{n-1,1},\ldots,a_{n-1,n-1}$ of the last row of the score matrix are mostly increasing, with most of the probability mass function being allocated to the last few coefficients, which contrasts to the behaviour of the attention mechanism on language related tasks [36]. Even though both NLP and time series forecasting deal with sequences of data, the main difference relies on the nature of this data. In NLP words are embedded into a vector space so that words related in meaning have similar dot product. Moreover, words with related meaning may come up far apart in a paragraph whereas consecutive words may have small dot product. On the other hand, most variables in time series problems are continuous and, thus, data tends to be extremely similar around a particular timestamp, hence leading to the larger attention scores in that region. This motivates the following working hypothesis for this research.
@@ 3 Local Attention Mechanism (LAM)
Previous attempts have exploited this sparsity properties of self-attention in time series modelling to design selective masking strategies that, for each row of the attention scores matrix, set most entries to zero. For example, the LogSparse Transformer [28] (denoted LogTrans in our experimentation) uses, for each key with timestamp $t_{0}$, attention scores with those queries whose timestamp $t$ is in a local neighbourhood of $t_{0}$ and those queries whose timestamp $t$ satisfies $t_{0}-t=2^{j}$ for some positive integer $j$. The LongFormer [37] extends the above work to a more complicated sparse configuration in the realm of NLP. These masking strategies partially resolve the loss of positional information problem of the attention mechanism as the input is only partially combined at each attention layer and, thus, Proposition 1 does not hold for such mechanisms. However, it is not clear how to obtain an improvement in time and memory complexity, as these attention mechanisms have not been written in tensor algebra in an efficient way. Instead, sparse matrices are used to perform computations and represent the attention matrix, leading to minor memory and computational efficiency improvements.
In this work, we propose the Local Attention Mechanism (LAM), which only computes dot products on nearby inputs of the attention layer. As we will see, our proposed attention can be implemented efficiently in tensor algebra, leading to almost linear time and memory complexity in practice.
@@
Note that these dot products are exactly the ones not masked by $M$ in the definition of LAM. In our algorithm, we compute $12L-$ 1 dot products for each row of $Q$ instead of the $L$ dot products described in Eq. (4). The extra number of dot products computed will allow us to write LAM in tensor algebra, without affecting the time and memory complexity. We formally introduce this “splitting” idea in Definition 2.
@@
We note that $(r-1)L+1+j_{1}<0$ only when $j_{1}<(1-r)L-1$, which only occurs in the case $r=0$ and $j_{1}\leq L-2$. Indeed, for $r=0$ in Eq. (4), we only have to compute the dot product $\langle Q_{i_{1}},K_{j}\rangle$ for $0\leq j\leq i_{1}$. The extra zeroes added in $T_{k}$ for $r=0$ allow us to maintain the dimension of $T_{K}[r,\cdot,\cdot]$ across all values of $r$. Lemma 2 allows us to go back and forth between the indexes $(i,j)$ and $(i_{1},j_{1})$ for the changes of variables $i=rL+i_{1}$ and $j=(r-1)L+1+j_{1}$.
@@
In view of Lemma 2, adding $T_{M}$ to $T_{A}$ acts as the same masking as adding $M$ to $QK^{T}$, we are just operating on the variables $(i_{1},j_{1})$ instead of $(i,j)$. Figure 3 visualises the computation $\operatorname{softmax}(T_{A}+T_{M})$.
@@
Proof.
Applying the operator $\operatorname{softmax}$ to the $i$-th row of $(QK^{T}+M)$, leads to all entries equal to zero except for those entries indexed by $(i,j)$ with $i-L+1\leq j\leq i$ (a total of $L$ entries). Analogously, applying the operator $\operatorname{softmax}$ to $(T_{A}+T_{M})[r,i_{1},\cdot]$, leads to all entries equal to zero except for those entries indexed by $(i_{1},j_{1})$ with $i_{1}\leq j_{1}\leq i_{1}+L-1$ (a total of $L$ entries). Moreover, $L$ non-zero entries are the same in both cases. Summarising, we have shown that:
  • when $i-L+1\leq j\leq i$, we have $S[i,j]=T_{S}[r,i_{1},j_{1}]$.
This dichotomy concludes the proof. ∎
@@
Lemma 5.
Proof.
[…]
which equals $LAM(Q,K,V)[i,\cdot]$, as we wanted to prove. ∎
@@
0:  $Q$ and $K$ have dimensions $n\times d_{q}$ and $V$ has dimensions $n\times d_{v}$. Let $s=n/L$. Let $T_{M}$ be the tensor with dimensions $s\times L\times(2L-1)$ such that $T_{M}[r,i_{1},j_{1}]=0$ when $i_{1}\leq j_{1}\leq i_{1}+L-1$ and $T_{M}[r,i_{1},j_{1}]=-\infty$ otherwise. Here the tensor $T_{M}$ is a local mask, that will be applied to each one of the $s$ ”splits" or blocks of $QK^{T}$ that we compute, emulating the work of the mask $M$ in Eq. (3). This tensor is pre-computed.
@@
4:  $T_{A}=T_{Q}T_{K}^{T}$. Note that $T_{A}$ has dimensions $s\times L\times(2L-1)$ and that the matrix multiplication is on the last two dimensions of the tensor.
5:  $T_{S}=\operatorname{SoftMax}(T_{A}+T_{M})$.
@@
We can now prove the main theoretical result of this work on combining the lemmas proved in this section and counting how many steps are performed in Algorithm 1.
@@
Proof.
This follows from instantiating $L$ in Theorem 6. ∎
4 Empirical comparison to the state of the art
In this section, we present the experimental setup and results yielded by the vanilla transformer fuelled by our attention mechanism, LAM, and compare these to the state of the art in time-series forecasting via transformers. In this section we follow the evaluation framework proposed in the Informer [22] paper. We will begin with a brief explanation of the datasets utilised and the algorithms involved in Section 4.1, moreover, we will provide details regarding the hardware used in our experiments and the open-access repository where the code is published. In Section 4.2, we analyse the results obtained including a Bayes test analysis. In Section 4.3 we perform a comparison between the attention mechanisms LAM and ProbAttn utilising the vanilla transformer architecture. Finally, in Section 4.4 we argue that the datasets that are used in the literature for benchmarking of models for the LSTF problem are not sufficient, which motivates Section 5.
@@ 4.1 Experimental Details
Datasets: For the sake of fairness we first follow the experimental methodology of Informer [22], we use four subsets of ETT: ETTh1 and ETTh2 (with hourly sampling), and ETTm1 and ETTm2 (with a 15-minute granularity). The ECL dataset records electricity consumption data for 321 clients, converted into hourly consumption, with ’MT 320’ as the target variable. The Weather dataset consists of climatological data from approximately 1,600 locations in the United States, collected hourly over four years, featuring ’wet bulb’ as the target variable along with 11 other climate features. Each dataset was divided into training, validation, and test sets. Details of the datasets can be observed in Table 1.
@@
Multivariate prediction, error metrics and validation: The primary objective of this comparison is to determine which model produces the least error in multivariate predictions on the datasets described above. Multivariate prediction involves using a fixed-size input sequence to predict a fixed-size output sequence across all variables in the dataset. The problem of multi-variable forecasting is significantly more challenging than single-variable forecasting. Incorporating more than one input variable allows for the extraction of more valuable information by capturing dependencies between variables. However, forecasting multiple variables simultaneously requires more sophisticated modelling, as errors can accumulate across each component. Furthermore, including more variables increases the model’s complexity, posing a challenge to assigning meaningful values to the weights. Consequently, the results obtained from multi-variable forecasting carry more significance and provide a more comprehensive evaluation of model performance. To ensure the validity and representativeness of the results, we have incorporated Bayesian tests [38], which serve to facilitate and formalise the evaluation process.
@@
Given the long training times needed in the LSTF problem, we present the results of one training execution, as it is standard in the literature. Due to the nature of time series, the methodology employed consists in dividing each dataset into training, validation, and test sets to monitor training and assess performance effectively. Specifically, the division is 12/4/4 months for ETT, 15/3/4 months for ECL, and 28/10/10 months for WTH as in [22].
Algorithms: We have maintained the same comparative framework as presented in the Informer paper, utilising the five algorithms selected by the original authors, being a representative sample of algorithms involving statistical methods and recurrent networks. This selection comprises ARIMA [39], Prophet [40], LSTMa [41], LST-Net [24], and DeepAR [42]. To better represent the state-of-the-art in time series forecasting, we have also included Reformer [43] and LogSparse self-attention [28], alongside the Informer model [22]. All these models have undergone hyperparameter optimisation to determine the most suitable configuration for each dataset.
@@
4.2 Local Attention Mechanism against the state-of-the-art
First, we will analyse a comparison between Informer and the number of parameters of the vanilla transformer incorporating LAM as attention mechanism, in order to test the improvement that arises from incorporating our proposal on vanilla architecture. It is important to highlight that the architecture accompanying our attention mechanism is exactly the vanilla architecture, based on Vaswani’s [20] original paper, without any modifications to the encoding-decoding framework or the positional encoding. The only difference between both models is the inclusion of LAM instead of the full attention mechanism. In the experimentation presented in this section the number of encoding and decoding layers in the vanilla transformer is 3, as opposed to 4 encoding layers and 2 decoding layers in the Informer architecture [22].
@@ 4.2 Local Attention Mechanism against the state-of-the-art
Table 2: Number of parameters of Informer and our Local Attention transformer model over the ECL dataset. The ratio is computed as Local Attention model parameters divided by Informer parameters. Seq_len parameter stands for the size of the input and prediction window. The number of encoding/decoding layers presents two numbers, $a/b$, where $a$ is the number of encoding layers in the model and $b$ is the number of decoding layers.
@@
Table 2 presents the size of both models on a fixed dataset while varying the size of the input window. Our model exhibits sensitivity to this variation, whereas Informer does not. The reason for this is that Informer has a fixed architecture that accepts an input window of size at most 512 time-steps. For smaller input window sizes, the input is extended with zeroes until reaching the mentioned size. Note that all the parameters in the Informer model are still used in the model independently of the size of the input window. In the optimal scenario, we achieve a parameter reduction factor of $2.1$, and in the worst scenario, a factor of $1.9$. Therefore, the model incorporating LAM is significantly more compact in terms of the number of parameters compared to Informer, and consequently, in terms of memory usage. This compactness could allow us to concatenate more layers of our attention mechanism, thereby more accurately capturing long-term temporal dependencies. However, to highlight the effect of our attention mechanism in prediction accuracy, we have chosen to have a similar number of encoding and decoding layers as Informer.
Methods LAM Informer LogTrans Reformer LSTMa LSTnet
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 24 0.471 0.448 0.577 0.549 0.686 0.604 0.991 0.754 0.650 0.624 1.293 0.901
48 0.560 0.545 0.685 0.625 0.766 0.757 1.313 0.906 0.702 0.675 1.456 0.960
168 1.011 0.807 0.931 0.752 1.002 0.846 1.824 1.138 1.212 0.867 1.997 1.214
336 0.923 0.756 1.128 0.873 1.362 0.952 2.117 1.280 1.424 0.994 2.655 1.369
720 0.993 0.732 1.215 0.896 1.397 1.291 2.415 1.520 1.960 1.322 2.143 1.380
ETTh2 24 0.588 0.543 0.720 0.665 0.828 0.750 1.531 1.613 1.143 0.813 2.742 1.457
48 0.780 0.692 1.457 1.001 1.806 1.034 1.871 1.735 1.671 1.221 3.567 1.687
168 3.209 1.479 3.489 1.515 4.070 1.681 4.660 1.846 4.117 1.674 3.242 2.513
336 6.289 2.192 2.723 1.340 3.875 1.763 4.028 1.688 3.434 1.549 2.544 2.591
720 2.835 1.204 3.467 1.473 3.913 1.552 5.381 2.015 3.963 1.788 4.625 3.709
ETTm1 24 0.264 0.301 0.323 0.369 0.419 0.412 0.724 0.607 0.621 0.629 1.968 1.170
48 0.714 0.561 0.494 0.503 0.507 0.583 1.098 0.777 1.392 0.939 1.999 1.215
96 0.487 0.511 0.678 0.614 0.768 0.792 1.433 0.945 1.339 0.913 2.762 1.542
288 0.522 0.522 1.056 0.786 1.462 1.320 1.820 1.094 1.740 1.124 1.257 2.076
672 0.774 0.658 1.192 0.926 1.669 1.461 2.187 1.232 2.736 1.555 1.917 2.941
Weather 24 0.307 0.349 0.335 0.381 0.435 0.477 0.655 0.583 0.546 0.570 0.615 0.545
48 0.409 0.450 0.395 0.459 0.426 0.495 0.729 0.666 0.829 0.677 0.660 0.589
168 0.527 0.537 0.608 0.567 0.727 0.671 1.318 0.855 1.038 0.835 0.748 0.647
336 0.582 0.576 0.702 0.620 0.754 0.670 1.930 1.167 1.657 1.059 0.782 0.683
720 0.762 0.670 0.831 0.731 0.885 0.773 2.726 1.575 1.536 1.109 0.851 0.757
ECL 48 0.294 0.379 0.344 0.393 0.355 0.418 1.404 0.999 0.486 0.572 0.369 0.445
168 0.282 0.375 0.368 0.424 0.368 0.432 1.515 1.069 0.574 0.602 0.394 0.476
336 0.287 0.375 0.381 0.431 0.373 0.439 1.601 1.104 0.886 0.795 0.419 0.477
720 0.368 0.391 0.406 0.443 0.409 0.454 2.009 1.170 1.676 1.095 0.556 0.565
960 0.417 0.497 0.460 0.548 0.477 0.589 2.141 1.387 1.591 1.128 0.605 0.599
Count 43 7 0 0 0 0
Table 3: Multivariate long sequence time-series forecasting results on the datasets described in Section 4.1.
@@
After comparing the sizes of the models, we proceed to examine the results of multivariate predictions. A comprehensive comparison of all algorithms for the ETTh1, ETTh2, ETTm1, Weather, and ECL datasets is presented in Table 3. The most notable aspect of the table is the superior performance of LAM. Out of 50 combinations of results and metrics, LAM ranks highest in 43, followed by Informer, which leads in 7. When compared to Reformer and LogTrans, our approach significantly outperforms their results. This superiority is even more pronounced in comparison to LSTMa and LST-Net, particularly with longer prediction horizons, where the performance gap is the largest. This can be attributed to the capability of our architecture to capture long-term dependencies. While LSTM-based algorithms perform well with short prediction horizons, LAM demonstrates substantial improvement and superior performance with longer prediction horizons, supporting the hypothesis that transformers are better equipped for LSTF problems.
@@
Focusing on a direct comparison between LAM and Informer allows for a more detailed case-by-case analysis. Overall, LAM outperforms Informer, particularly for large forecast horizons in the ETT datasets. For example, in the case of ETTh2 with a forecast window of $720$, LAM achieves an MSE of $2.8$ compared to Informer’s $3.4$. Although LAM also demonstrates superior performance in the ECL and Weather datasets, the margin of superiority is not as pronounced as in the ETT datasets.
(a) MSE Bayesian test
(b) MAE Bayesian test
Figure 5: Bayesian tests for Informer vs LAM
@@
To better support this comparison, we conducted two Bayesian tests: one focusing on the MSE and the other on the MAE, using a $0.05$ ROPE. As illustrated in Figure 5(a) and Figure 5(b), there is statistically significant superiority of LAM with respect to the rest of algorithms considered. These significant differences underscore the robustness and effectiveness of our model compared to existing LSTF algorithms. Consequently, this provides strong validation for our experimentation methodology and supports the reliability of the conclusions drawn from our study.
4.3 Comparing LAM and ProbAttn within the same architecture framework

### v3

@@ 2.4 The Encoder-Decoder architecture and Positional Encoding
3 Local Attention Mechanism
In this section we introduce LAM and an efficient algorithm, based on expressing LAM in tensor algebra, that leads to $\Theta(n\log n)$ memory and time complexity. This algorithm has been implemented in stock Pytorch and TensorFlow without the need for any low level CUDA code.
3.1 Theoretical definition
The high-level fundamental behind LAM is exploiting locality in order to re-define the attention mechanism. Previous studies on time series have empirically shown that the rows of the attention scores matrix behave as long tail distributions (li2019enhancing, ; zhou2023informer, ). This behaviour is present at any layer independently of the depth. For example, the coefficients $a_{n-1,0},a_{n-1,1},\ldots,a_{n-1,n-1}$ of the last row of the score matrix are mostly increasing, with most of the probability mass function being allocated to the last few coefficients, which contrasts to the behaviour of the attention mechanism on language related tasks (galassi2020attention, ). Even though both NLP and time series forecasting deal with sequences of data, the main difference relies on the nature of this data. In NLP words are embedded into a vector space so that words related in meaning have similar dot product. Moreover, words with related meaning may come up far apart in a paragraph whereas consecutive words may have small dot product. On the other hand, most variables in time series problems are continuous and, thus, data tends to be extremely similar around a particular timestamp, hence leading to the larger attention scores in that region. This motivates the following working hypothesis for this research.
@@ 3.1 Theoretical definition
Previous attempts have exploited this sparsity properties of self-attention in time series modelling to design selective masking strategies that, for each row of the attention scores matrix, set most entries to zero. For example, the LogSparse Transformer (li2019enhancing, ) (denoted LogTrans in our experimentation) uses, for each key with timestamp $t_{0}$, attention scores with those queries whose timestamp $t$ is in a local neighbourhood of $t_{0}$ and those queries whose timestamp $t$ satisfies $t_{0}-t=2^{j}$ for some positive integer $j$. The LongFormer (beltagy2020longformer, ) extends the above work to a more complicated sparse configuration in the realm of NLP. These masking strategies partially resolve the loss of positional information problem of the attention mechanism as the input is only partially combined at each attention layer and, thus, Proposition 1 does not hold for such mechanisms. However, it is not clear how to obtain an improvement in time and memory complexity, as these attention mechanisms have not been written in tensor algebra in an efficient way. Instead, sparse matrices are used to perform computations and represent the attention matrix, leading to minor memory and computational efficiency improvements.
In this work, we propose the Local Attention Mechanism (LAM), which only computes dot products on nearby inputs of the attention layer. As we will see, LAM can be implemented efficiently in tensor algebra, leading to almost linear time and memory complexity in practice.
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 LAM implementation
[…]
@@
[…]
@@
We can now provide the main theoretical result of this work on combining the lemmas proved in this section and counting how many steps are performed in Algorithm 1. All proofs from this section are available in Appendix A.
@@
4 LAM empirical comparison to the state of the art
In this section, we present the experimental setup and results yielded by the vanilla transformer, fuelled by LAM, and compare these to the state of the art in time-series forecasting via transformers. In this section we follow the evaluation framework proposed in the Informer (zhou2023informer, ) paper. We will begin with a brief explanation of the datasets utilised and the algorithms involved in Section 4.1, moreover, we will provide details regarding the hardware used in our experiments and the open-access repository where the code is published. In Section 4.2, we analyse the results obtained including a Bayes test analysis. In Section 5 we perform a comparison between the attention mechanisms LAM and ProbAttn utilising the vanilla transformer architecture. Finally, in Section 5.2 we argue that the datasets that are used in the literature for benchmarking of models for the LSTF problem are not sufficient, which motivates Section 6.
@@ 4.1 Experimental Details
Datasets: For the sake of fairness we first follow the experimental methodology of Informer (zhou2023informer, ), we use four subsets of ETT: ETTh1 and ETTh2 (with hourly sampling), and ETTm1 and ETTm2 (with a 15-minute granularity). The ECL dataset records electricity consumption data for 321 clients, converted into hourly consumption, with ’MT 320’ as the target variable. The Weather dataset consists of climatological data from approximately 1,600 locations in the United States, collected hourly over four years, featuring ’wet bulb’ as the target variable along with 11 other climate features. Each dataset was divided into training, validation, and test sets. Details of the datasets can be observed in Table 1.
@@
Multivariate prediction, error metrics and validation: The primary objective of this comparison is to determine which model produces the least error in multivariate predictions on the datasets described above. Multivariate prediction involves using a fixed-size input sequence to predict a fixed-size output sequence across all variables in the dataset. The problem of multi-variable forecasting is significantly more challenging than single-variable forecasting. Incorporating more than one input variable allows for the extraction of more valuable information by capturing dependencies between variables. However, forecasting multiple variables simultaneously requires more sophisticated modelling, as errors can accumulate across each component. Furthermore, including more variables increases the model’s complexity, posing a challenge to assigning meaningful values to the weights. Consequently, the results obtained from multi-variable forecasting carry more significance and provide a more comprehensive evaluation of model performance. To ensure the validity and representativeness of the results, we have incorporated Bayesian tests (benavoli2014bayesian, ), which serve to facilitate and formalise the evaluation process.
@@
Given the long training times needed in the LSTF problem, we present the results of one training execution, as it is standard in the literature. Due to the nature of time series, the methodology employed consists in dividing each dataset into training, validation, and test sets to monitor training and assess performance effectively. Specifically, the division is 12/4/4 months for ETT, 15/3/4 months for ECL, and 28/10/10 months for WTH as in (zhou2023informer, ).
Algorithms: We have maintained the same comparative framework as presented in the Informer paper, utilising the five algorithms selected by the original authors, being a representative sample of algorithms involving statistical methods and recurrent networks. This selection comprises ARIMA (ariyo2014stock, ), Prophet (taylor2018forecasting, ), LSTMa (bahdanau2014neural, ), LST-Net (lai2018modeling, ), and DeepAR (salinas2020deepar, ). To better represent the state-of-the-art in time series forecasting, we have also included Reformer (kitaev2020reformer, ), Autoformer (wu2021autoformer, ) and LogSparse self-attention (li2019enhancing, ), alongside the Informer model (zhou2023informer, ). All these models have undergone hyperparameter optimisation to determine the most suitable configuration for each dataset.
@@
4.2 Local Attention Mechanism compared to the state-of-the-art
First, we will analyse a comparison between Informer and the number of parameters of the vanilla transformer incorporating LAM as attention mechanism, in order to test the improvement that arises from incorporating LAM on vanilla architecture. It is important to highlight that the architecture accompanying our attention mechanism is exactly the vanilla architecture, based on Vaswani’s (vaswani2017attention, ) original paper, without any modifications to the encoding-decoding framework or the positional encoding. The only difference between both models is the inclusion of LAM instead of the full attention mechanism. In the experimentation presented in this section the number of encoding and decoding layers in the vanilla transformer is 3, as opposed to 4 encoding layers and 2 decoding layers in the Informer architecture (zhou2023informer, ).
@@ 4.2 Local Attention Mechanism compared to the state-of-the-art
Table 2: Number of parameters for Informer and LAM model over the ECL dataset. The ratio is computed as LAM model parameters divided by Informer parameters.
@@
Table 2 presents the size of both models on a fixed dataset while varying the size of the input window. Our model exhibits sensitivity to this variation, whereas Informer does not. The reason for this is that Informer has a fixed architecture that accepts an input window of size at most 512 time-steps. For smaller input window sizes, the input is extended with zeroes until reaching the mentioned size. Note that all the parameters in the Informer model are still used in the model independently of the size of the input window. In the optimal scenario, we achieve a parameter reduction factor of 2.12.12.1, and in the worst scenario, a factor of 1.91.91.9. Therefore, the model incorporating LAM is significantly more compact in terms of the number of parameters compared to Informer, and consequently, in terms of memory usage. This compactness could allow us to concatenate more layers of our attention mechanism, thereby more accurately capturing long-term temporal dependencies. However, to highlight the effect of our attention mechanism in prediction accuracy, we have chosen to have a similar number of encoding and decoding layers as Informer.
Methods LAM Informer LogTrans Reformer Autoformer LSTMa LSTnet
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
24 0.471 0.448 0.577 0.549 0.686 0.604 0.991 0.754 0.672 0.548 0.650 0.624 1.293 0.901
48 0.560 0.545 0.685 0.625 0.766 0.757 1.313 0.906 0.802 0.606 0.702 0.675 1.456 0.960
168 1.011 0.807 0.931 0.752 1.002 0.846 1.824 1.138 1.003 0.684 1.212 0.867 1.997 1.214
336 0.923 0.756 1.128 0.873 1.362 0.952 2.117 1.280 1.183 0.787 1.424 0.994 2.655 1.369
720 0.993 0.732 1.215 0.896 1.397 1.291 2.415 1.520 1.254 0.853 1.960 1.322 2.143 1.380
ETTh2
24 0.588 0.543 0.720 0.665 0.828 0.750 1.531 1.613 0.923 0.998 1.143 0.813 2.742 1.457
48 0.780 0.692 1.457 1.001 1.806 1.034 1.871 1.735 1.530 1.207 1.671 1.221 3.567 1.687
168 3.209 1.479 3.489 1.515 4.070 1.681 4.660 1.846 3.956 1.958 4.117 1.674 3.242 2.513
336 6.289 2.192 2.723 1.340 3.875 1.763 4.028 1.688 4.016 1.897 3.434 1.549 2.544 2.591
720 2.835 1.204 3.467 1.473 3.913 1.552 5.381 2.015 4.235 2.004 3.963 1.788 4.625 3.709
ETTm1
24 0.264 0.301 0.323 0.369 0.419 0.412 0.724 0.607 0.867 0.594 0.621 0.629 1.968 1.170
48 0.714 0.561 0.494 0.503 0.507 0.583 1.098 0.777 1.093 0.692 1.392 0.939 1.999 1.215
96 0.487 0.511 0.678 0.614 0.768 0.792 1.433 0.945 0.746 0.583 1.339 0.913 2.762 1.542
288 0.522 0.522 1.056 0.786 1.462 1.320 1.820 1.094 1.007 0.697 1.740 1.124 1.257 2.076
672 0.774 0.658 1.192 0.926 1.669 1.461 2.187 1.232 1.119 0.771 2.736 1.555 1.917 2.941
Weather
24 0.307 0.349 0.335 0.381 0.435 0.477 0.655 0.583 0.564 0.502 0.546 0.570 0.615 0.545
48 0.409 0.450 0.395 0.459 0.426 0.495 0.729 0.666 0.588 0.544 0.829 0.677 0.660 0.589
168 0.527 0.537 0.608 0.567 0.727 0.671 1.318 0.855 0.716 0.637 1.038 0.835 0.748 0.647
336 0.582 0.576 0.702 0.620 0.754 0.670 1.930 1.167 0.771 0.678 1.657 1.059 0.782 0.683
720 0.762 0.670 0.831 0.731 0.885 0.773 2.726 1.575 0.848 0.761 1.536 1.109 0.851 0.757
ECL
48 0.294 0.379 0.344 0.393 0.355 0.418 1.404 0.999 0.352 0.459 0.486 0.572 0.369 0.445
168 0.282 0.375 0.368 0.424 0.368 0.432 1.515 1.069 0.397 0.492 0.574 0.602 0.394 0.476
336 0.287 0.375 0.381 0.431 0.373 0.439 1.601 1.104 0.456 0.547 0.886 0.795 0.419 0.477
720 0.368 0.391 0.406 0.443 0.409 0.454 2.009 1.170 0.46 0.529 1.676 1.095 0.556 0.565
960 0.417 0.497 0.460 0.548 0.477 0.589 2.141 1.387 0.601 0.610 1.591 1.128 0.605 0.599
Count 43 7 0 0 0 0 0
Table 3: Multivariate LSTF results on the datasets described in Section 4.1.
@@
After comparing the sizes of the models, we proceed to examine the results of multivariate predictions. A comprehensive comparison of all algorithms for the ETTh1, ETTh2, ETTm1, Weather, and ECL datasets is presented in Table 3. The most notable aspect of the table is the superior performance of LAM. Out of 50 combinations of results and metrics, LAM ranks highest in 43, followed by Informer, which leads in 7. When compared to Reformer, Autoformer and LogTrans, our approach significantly outperforms their results. This superiority is even more pronounced in comparison to LSTMa and LST-Net, particularly with longer prediction horizons, where the performance gap is the largest. This can be attributed to the capability of our architecture to capture long-term dependencies. While LSTM-based algorithms perform well with short prediction horizons, LAM demonstrates substantial improvement and superior performance with longer prediction horizons, supporting the hypothesis that transformers are better equipped for LSTF problems.
@@
Focusing on a direct comparison between LAM and Informer allows for a more detailed case-by-case analysis. Overall, LAM outperforms Informer, particularly for large forecast horizons in the ETT datasets. For example, in the case of ETTh2 with a forecast window of 720720720, LAM achieves an MSE of 2.82.82.8 compared to Informer’s 3.43.43.4. Although LAM also demonstrates superior performance in the ECL and Weather datasets, the margin of superiority is not as pronounced as in the ETT datasets.
(a) MSE Bayesian test.
(b) MAE Bayesian test.
Figure 5: Bayesian tests for Informer vs LAM.
@@
To better support this comparison, we conducted two Bayesian tests: one focusing on the MSE and the other on the MAE, using a 0.050.050.05 ROPE. As illustrated in Figure 5(a) and Figure 5(b), there is statistically significant superiority of LAM with respect to the rest of algorithms considered. These significant differences underscore the robustness and effectiveness of our model compared to existing LSTF algorithms. Consequently, this provides strong validation for our experimentation methodology and supports the reliability of the conclusions drawn from our study.
5 Comparing LAM and ProbAttn within the same architecture framework
Following the previous comparison, it is pertinent to consider whether the two attention mechanisms, LAM and Informer’s attention mechanism (ProbAttn) are competing on equal terms. In this Section we will conduct an experimental comparison between ProbAttention and LAM within the transformer vanilla architecture. The same validation schema from the previous section is maintained. The results derived from this section will help us have an empirical comparison focusing exclusively on attention mechanisms.
5.1 Experimental results